Hello again, I wonder if you can help me with these word problems?
Suppose that x liters of a 31% acid solution are mixed with y liters of a 44% solution to obtain 100 L of a 36% solution. One equation in a system is x+y=100 what is the second equation?

Bay City and Vernonville are 176 miles apart. A car leaves Bay City traveling towards Vernonville, and another car leaves Vernonville at the same time, traveling towards Bay City. The car leaving Bay City averages 10 miles per hour more than the other, and they meet after 1 hour and 36 minutes. What are the average speeds of the cars?
I got: Bay City 60 and Vernonville 50
I really need these by the 14th!!! Thank you fro all of your help!

Answer by Alan3354(69443)

Suppose that x liters of a 31% acid solution are mixed with y liters of a 44% solution to obtain 100 L of a 36% solution.
One equation in a system is x+y=100 what is the second equation?
x + y = 100 (total volume)
0.31x + 0.44y = 0.36*100 (amount of acid)
----------------------------
Bay City and Vernonville are 176 miles apart. A car leaves Bay City traveling towards Vernonville, and another car leaves Vernonville at the same time, traveling towards Bay City. The car leaving Bay City averages 10 miles per hour more than the other, and they meet after 1 hour and 36 minutes. What are the average speeds of the cars?
I got: Bay City 60 and Vernonville 50
The speeds are added when going in opposite directions.
They travel 176 miles in 1.6 hours = 110 mph
50 and 60 is correct.